OpenStudy (anonymous):

A police car at rest is passed by a speeder traveling at a constant speed of 30m/s. the police takes off immediately in pursuit of the speeder and catches up to the speeder in 750m maintaining a constant acceleration. A). How long did it take the police car to catch up with the speeder? B). What was the acceleration of the police car? C). What is the speed of the car when it overtakes the speeder?

OpenStudy (anonymous):

alright, these questions are best solved using the kinematic equations, there are 3 of them, are you familiar with them?

OpenStudy (anonymous):

@dwade you around?

OpenStudy (anonymous):

@DemolisionWolf not quite, I haven't memorized them :3

OpenStudy (anonymous):

hey no worries, as long as you know what they even are! this first part is a bit tricky... and looks like this \[\frac{ 750 meters }{ 1 } \times \frac{ seconds }{ 30 meters }\] the idea is that we know the cop catches up to the car in 750 meters, and what that is really saying is the cop goes from 0m/s to 30m/s over a length of 750 meters. so we set up the fractions as seen. then multiply them to get the answer of how many seconds

OpenStudy (anonymous):

so you end up getting 25 seconds?

OpenStudy (anonymous):

yep, now part B...

OpenStudy (anonymous):

we will use 1 of the 3 kinematic equations. looking over what we know, and what we are looking for, any of the three we can use, but this one is the easiest because it has the fewest variables. V = V + at the left V is the final velocity (meters/ second) the right V is the starting veloicty (meters/ second the a is for acceleration (meters / second^2) the t is for time, (seconds) plug in the values you know, then solve for a.

OpenStudy (anonymous):

@DemolisionWolf would the equation read 30m/s = 0m/s + (30m/s^2)(25second)

OpenStudy (anonymous):

correct, except leave in the a, 30m/s = 0m/s + (a)(25second) becuase we are solving for a, it is unknown at this point. so can you solve for a in the equation?

OpenStudy (anonymous):

a = 5 m/s?

OpenStudy (anonymous):

not quite, i think you have the right idea. lets just work on your algebra 30m/s = 0m/s + (a)(25second) when solving for a variable, we want to move all the terms that don't have a variable attached to them to one side, and all the terms with a variable attached to them to the other side of the = sign. it's called 'isolating the variable' for us, we already have all the terms that are just numbers to one side, and left all the terms with a variable to the other side. 30 = (a)(25) so we will divide both sides by 25, inorder to isolate a 30/25 = a

OpenStudy (anonymous):

oh yeah i was a little hesitant on that so i get 1.2 m/s

OpenStudy (anonymous):

a = 1.2

OpenStudy (anonymous):

@DemolisionWolf you there?

OpenStudy (anonymous):

hey sorry, i'm making bread today, so i'm in and out of the kitchen!

OpenStudy (anonymous):

yes a = 1.2 m/s^2 is correct

OpenStudy (anonymous):

as for part C, as far as I can tell, there is no equations needed, just logic

OpenStudy (anonymous):

oh man im sorry take your time i know when my mother makes bread she's focused on it. cant neglect it.

OpenStudy (anonymous):

alright sounds good ill work it out

OpenStudy (anonymous):

so C is goign to be the same speed as the car the cop is chasing